# OrionNav: Online Planning for Robot Autonomy with Context-Aware LLM and Open-Vocabulary Semantic Scene Graphs

### v1

@@ III Problem Formulation
We address the challenge of developing an autonomous language-driven physical agent capable of operating in dynamic, unknown environments. The agent, a quadruped robot, is equipped with an RGB-D camera, LiDAR, and leg odometry sensors (IMU, joint encoders, and contact sensors). The primary objective is to enable the robot to autonomously search for and navigate to specific objects or rooms in response to natural language queries issued by a remote operator. The robot must operate without prior knowledge of the environment, which consists of multiple interconnected rooms containing various objects. The environment is subject to change over time, with no predetermined types of objects, allowing for arbitrary categories and adding complexity to the search and navigation tasks. The robot must interpret queries, explore the environment, and locate targets in real time while adapting its strategies when initial attempts are unsuccessful. Performance is evaluated based on the robot’s ability to successfully locate and reach queried targets, considering time and path efficiency. The problem encompasses challenges in perception, navigation, natural language understanding, and decision-making in unstructured, real-world settings. The desired outcome is a system that can build and maintain a scalable representation of the environment and objects within it, which can be utilized to handle subsequent queries more efficiently.
@@ IV METHODOLOGY
This section presents our novel framework for addressing the challenges outlined in Section III. Our framework efficiently explores the environment, constructing a hierarchical scene graph that captures object semantics and relationships (Section IV-A1). An LLM-based high-level planner translates human queries into structured commands for execution using the scene graph (Section IV-B). The robot employs safe motion primitives for exploration and navigation, reducing accident risk (Section IV-C). This methodology integrates open-vocabulary semantic mapping, hierarchical scene graph generation, and LLM-based query interpretation with a low-level motion planner and controller, enabling safe autonomous operation in response to the problem outlined in Section III. These modules, as detailed in the following subsections, can be easily substituted as advancements occur in these areas or to accommodate different user needs, provided they adhere to the communication protocols.
@@ IV-A1 Open-vocabulary Semantic Segmentation
The agent uses RGB frames from the RGB-D camera to semantically segment the scene. We use an off-the-shelf open-vocabulary semantic segmentation framework, FC-CLIP [93], that comprises of a class agnostic mask generator, in-vocabulary classifier, and out-of-vocabulary classifier. All modules leverage multi-scale feature maps extracted from a frozen ConvNeXt-L CLIP (Contrastive Language-Image Pretraining) backbone [94]. The mask generator passes these feature maps through a pixel decoder with multi-scale deformable attention, as described in Mask2Former [98], to enhance the pixel features. These enhanced features, along with object query embeddings, are then passed through a series of mask decoders. Each decoder consists of masked cross-attention, self-attention, and a feed-forward network.
@@
Figure 3: Prompt templates used to generate text embeddings for each category. In these templates, every instance of ‘<label>’ is replaced with corresponding category names.
@@
The candidate masks in FC-CLIP require category text embeddings for each class to enable zero-shot classification. These embeddings are generated using a method similar to CLIP, employing a set of predefined template prompts (e.g., ‘a photo of a <label>’, ‘an image of a <label>’) for each class as shown in Fig. 3. For each class, the class name and its synonyms (e.g., cabinet, cupboard, bureau, china cabinet) are individually inserted into these templates, and the resulting texts are processed through a CLIP text encoder. Importantly, each synonym generates its own unique embedding, allowing the model to capture subtle semantic differences. However, during the classification process, these separate embeddings are all associated with a single, general class name (e.g., ‘cabinet’) for output, which aids in LLM room classification and scene graph generation. This approach allows for on-the-fly computation of embeddings for new objects without additional training and enhances the model’s robustness to different object naming conventions. The classification process involves both out-of-vocabulary and in-vocabulary classifiers. The out-of-vocabulary classifier, which applies mask pooling to the frozen CLIP backbone’s feature maps, is designed to maintain CLIP’s open-vocabulary recognition ability for unseen classes. The in-vocabulary classifier, which pools the pixel decoder’s output feature maps, is optimized for better performance on classes seen during training. Both compute classification scores using cosine similarity between the pooled features and each of the synonym embeddings, followed by softmax with a temperature of 0.07. The highest score among synonyms determines the class prediction. These scores from both classifiers are then combined using geometric ensembling, leveraging the strengths of both classifiers to improve overall performance on both seen and unseen classes. Importantly, experiments have shown that using both classifiers also improves performance on out-of-vocabulary classes, demonstrating the synergistic effect of this approach. The mask generator produces 100 fixed mask proposals per image, which are filtered based on classification scores and background class assignment. Remaining masks $\mathcal{M}_{o}$ are associated with their corresponding object class $o$, using the general class name. Objects like ceilings, floors, and walls are segmented but excluded from object mapping to improve robustness against outlier masks. Category text embeddings for COCO Panoptic dataset classes (used to train FC-CLIP) and additional labels relevant to indoor environments, along with their synonyms, are included in the experiments, all mapping to appropriate general class names. This process enables FC-CLIP to perform effective zero-shot segmentation across a wide range of object classes, including those not seen during training, while being flexible to various object naming conventions and providing consistent class names for downstream tasks.
@@ IV-A2 Semantic Object Mapping
The predicted 2D segmentation mask for an object, as described in Section IV-A1 and denoted by $\mathcal{M}_{o}$ (for object $o$) is used to estimate the object’s center $(u_{o},v_{o})$ in the camera frame. These center points are then projected from the camera plane to 3D coordinates $P_{o}^{c}=(X_{o}^{c},Y_{o}^{c},Z_{o}^{c})$ in the camera frame, using depth frame from RGB-D camera, and camera intrinsics. This approach was selected over alternatives such as using 3D bounding boxes or computing the mean of the pointcloud, considering simplicity and computational complexity. These object observations are subsequently converted from the camera frame to the world frame using the pose from the robot’s odometry. We employ an off-the-shelf SLAM toolbox [99] that utilizes the robot’s onboard LiDAR and legged-inertial odometry to build a 2D map of the environment as well as estimate the pose of the robot. The detected objects ($P_{o}^{c}$ in the camera frame) are transformed to the world frame $P_{o}^{w}$ using the estimated pose and associated with previous map objects.
@@
The map objects are ones observed in prior frames and frame objects are objects observed in the current frame. All frame objects observed in the first frame are considered as map objects and if there are multiple map objects belonging to the same class, they are given a different instance number. The instance number is the instance count of a particular class in the scene and is thus initialized at 1 and incremented with new observations. For the subsequent frames, the frame object and map object are associated using density-based clustering (DB-SCAN) [100] based assignment as it offers a good balance between accuracy and computational speed compared to more complex methods like Hungarian matching [101]. In this process, frame objects are projected into the world coordinate frame and DB-SCAN is applied to group frame and map objects of the same class, using a local radius threshold of $\epsilon$ meters. For each resulting cluster, only one map object is retained, while the others are discarded. DBSCAN is effectively able to handle grouping multiple nearby detections, such as when a large object is detected as several smaller bounding boxes across consecutive frames. We consider three cases for grouping objects: for larger objects like tables, doors, and noticeboards, we set $\epsilon=2$ meters to account for their size; for smaller objects, such as books, a reduced threshold of $\epsilon=0.5$ meters is used; and for all other objects, a threshold of $\epsilon=1$ meter is applied. This makes it a good choice to filter out extra bounding boxes corresponding to a single object without using any complex heuristics unlike other filtering methods. All the objects without any associations are assigned as new map objects with unique instance numbers, thus forming a semantic object map $\mathcal{S}$.
@@
The system handles updates in object locations following loop closure and local bundle adjustment using a pose graph-based approach. Each object in the semantic object map $\mathcal{S}$ is associated with the keyframe in which it was first observed. A relative transform $T_{o}^{k}$ between the object and its associated keyframe in the pose graph is maintained. When loop closure or local bundle adjustment occurs and keyframe poses are updated, the location of each object is adjusted by applying the updated keyframe pose to the stored relative transform: $T_{o}^{w}=T_{k}^{w}\cdot T_{o}^{k}$, where $T_{o}^{w}$ is the updated world pose of the object, $T_{k}^{w}$ is the updated pose of the keyframe in the world coordinates, and $T_{o}^{k}$ is the relative transform between the object and its keyframe. An alternative approach involving bundle adjustment using all keyframes that observe an object was considered but not chosen due to its higher complexity and computational cost, favoring simplicity and efficiency in our current system. Our method ensures that object locations remain consistent with the refined map and robot trajectory, maintaining the integrity of the semantic object map even as the underlying SLAM system refines its estimates.
@@ IV-A3 Hierarchical Scene Graph
Indoor environments are, by design, distributed into several semantic components like floors and rooms. Based on the purpose and the objects present in these areas, rooms can be labeled into commonly known language terms. Since these terms are fairly common and consistent across all indoor environments, they become highly informative semantic labels. For e.g., a room labeled as a ‘kitchen’ would have a much higher probability of containing a microwave as opposed to another room labeled as an ‘office’. Hence, as proposed and utilized in previous works [44, 45, 15, 16], a 3D scene graph with room hierarchy is an effective and efficient representation of indoor spaces.
Typically most indoor spaces consist of several rooms connected by walkways called corridors. Hence, it is natural to find multiple clusters of objects while exploring such environments. Making use of this observed property, we propose the classification of the scene into multiple areas, obtained by applying density-based clustering (DB-SCAN [100]) on the coordinates of the map objects. While most previous works rely on language-aligned visual embedding models to find the best match, our experimentation has led us to believe that it is fairly difficult to model the influence of certain combinations of objects toward accurate room classification. Hence, we harness the capabilities of LLMs to generate responses that are consistent with the general notion of such semantic concepts. Providing the labels of all the objects present in each area cluster, we query an LLM to realize the most appropriate room label. To maintain consistency across the environment, we also provide a list of candidate room labels and a few general guidelines in the prompts. To generate these labels in real-time and on-device, we use a memory-efficient 4-bit quantized version of the LLama3 [102] model.
Finally, a tree structure is formed with each node containing properties such as the hierarchy level, text label, location in the map, and parents and children to store the scene graph efficiently (Fig. 2). This type of structure is easily converted to a Python dictionary format that can be directly passed to the LLM-based planner as a prompt.
@@ IV-B LLM Planner
12:           task_accomplished $\leftarrow$ ask_user(‘Is task done?’)
@@
The LLM Planner uses three action primitives to navigate and interact with the environment: goto(<room name>or <object name>), which navigates the robot to the specified room or near the object and then queries a visual agent to verify if the task is completed; search_room(<room name>), which first navigates the robot to the room given in the function argument and then rotates to detect nearby objects; and explore_globally(), which initiates a frontier-based exploration method. Documentation of these commands, including explanations and the scenarios in which they should be used, is provided to the LLM (see Fig. 4 for prompting details). To ground the LLM in the environment, we supply it with the hierarchical scene graph that stores information about the rooms and the objects within them, generated in parallel by a separate process and provided in JSON format. The LLM selects an action and fills in the arguments using the scene graph to accomplish the user-defined task, while also providing reasoning for its decision. It is also required to provide reasoning for it. The output is directly executed as a function call to the robot API, implemented using the Safe Robot Controller described in Section IV-C.
The plan generation framework operates as follows: If the scene graph is empty, the robot first performs a 360-degree rotation to observe the environment. The LLM is then tasked with generating an action based on the given task and the hierarchical scene graph. After this action is executed, if the task is not yet completed, the updated scene graph is provided, and the LLM is prompted to generate a subsequent action. In case an error occurs—due to incorrect syntax, low-level controller issues, or environmental conditions such as an inaccessible target point—a feedback message explaining the error is generated and provided to the LLM, prompting it to produce another action. Additionally, after the initial action is generated, a command history containing previously executed actions is supplied to prevent the repetition of the same actions. This iterative action generation process is formalized in Algorithm 1. The details of the action primitives provided to the LLM are as follows:
@@
This is the navigation primitive used by the LLM to move between rooms and approach objects. It can be used either to go near an object or navigate to a specific room. In the prompts, the Planner is instructed to specify an object name as well as the room it belongs to in the argument when the given task involves navigating to an object. If it is related to navigating to a specific room—like an office with three computers—only the room name is given in the arguments.
This command also allows the LLM Planner to determine whether the task has been completed. After navigating to the object or room specified in the arguments, it prompts a visual agent to verify whether the task has been accomplished. To prevent the Planner from overusing this command after every action, the LLM is conditioned to use it only when it believes the task is complete following the execution of this command. In the proposed framework, user feedback serves as the oracle visual agent. For a fully autonomous system capable of independently determining task completion, a Vision-Language Model (VLM) can be employed.
@@
This action navigates the robot to the specified room and then performs a 360-degree rotation there. It is designed to be used to explore a partially observed room, allowing the robot to detect objects within and update the room’s graph representation. While this command might be unnecessary in simulated environments with ground truth segmentation, it is crucial in our problem setting. Even if the robot has previously been in the room, imperfect perception may cause it to miss some objects. Therefore, this command provides the Planner with the opportunity to re-observe the room.
@@
When the scene graph is either empty or lacks sufficient information for the LLM to use the other commands,this primitive is expected to be utilized. It triggers a minute of frontier-based exploration (see Section IV-C3 for details), aiming to map new areas so that the LLM can provide more task-specific commands. After the exploration concludes, the robot performs a 360-degree rotation.
IV-C Safe Robotics Controller

### v2

@@ III Problem Formulation
We address the challenge of developing an autonomous language-driven physical agent capable of operating in dynamic, unknown environments. The agent, a quadruped robot, is equipped with an RGB-D camera, LiDAR, and leg odometry sensors (IMU, joint encoders, and contact sensors). The primary objective is to enable the robot to autonomously search for and navigate to specific objects or rooms in response to natural language queries issued by a remote operator. The robot must operate without prior knowledge of the environment, which consists of multiple interconnected rooms containing various objects. The environment is subject to change over time, with no predetermined types of objects, allowing for arbitrary categories and adding complexity to the search and navigation tasks. The robot must interpret queries, explore the environment, and locate targets in real time while adapting its strategies when initial attempts are unsuccessful. Performance is evaluated based on the robot’s ability to successfully locate and reach queried targets, considering time and path efficiency. The problem encompasses challenges in perception, navigation, natural language understanding, and decision-making in unstructured, real-world settings. The desired outcome is a system that can build and maintain a memory-efficient representation of the environment and objects within it, which can be utilized to handle subsequent queries more efficiently.
@@ IV Methodology
This section introduces OrionNav, our novel framework designed to address the challenges outlined in Section III. The framework efficiently explores the environment, constructing a hierarchical scene graph that captures object semantics and relationships (Section IV-A). To interpret human queries, OrionNav employs the LLM-based high-level planner that translates natural language into structured commands, which are then executed using the scene graph information (Section IV-B). It employs safe motion primitives for exploration and navigation, reducing accident risk (Section IV-C). By combining open-vocabulary semantic mapping, hierarchical scene graph generation, and LLM-based query interpretation with a low-level motion planner and controller, OrionNav enables safe autonomous operations that directly address the problem outlined in Section III. Importantly, these modules, which are further elaborated in the following subsections, are designed with modularity in mind, allowing for easy substitution as technological advancements occur or to meet varying user requirements, provided that new components adhere to the established communication protocols.
@@ IV-A1 Open-Vocabulary Semantic Segmentation
The agent uses color frames from the RGB-D camera to semantically segment the scene. We use an off-the-shelf open-vocabulary semantic segmentation framework, FC-CLIP [92], that comprises a class agnostic mask generator, in-vocabulary classifier, and out-of-vocabulary classifier. All modules leverage multi-scale feature maps extracted from a frozen ConvNeXt-L [93] CLIP [40] backbone. The mask generator passes these feature maps through a pixel decoder with multi-scale deformable attention, as described in Mask2Former [96], to enhance the pixel features. These enhanced features, along with object query embeddings, are then passed through a mask decoder with each layer in the decoder consisting of a masked cross-attention layer, a self-attention layer, and a feed-forward network. The object query embeddings are a fixed number (250) of learnable embeddings representing each potential object the model can predict. The mask decoder predicts a fixed set of candidate masks (100 masks) representing the segmentation of each object in the scene. The number of object queries and masks are independent, allowing flexible prediction of varying object counts while improving computational efficiency by generating masks only for valid predictions.
@@
Figure 3: Prompt templates used to generate text embeddings for each category. Each instance of ‘<label>’ is replaced with corresponding category names.
@@
The candidate masks in FC-CLIP require category text embeddings for each class to enable zero-shot classification. These embeddings of dimension 768, are generated using a method similar to CLIP, employing a set of predefined template prompts (e.g., ‘a photo of a <label>’, ‘an image of a <label>’) for each class as shown in Fig. 3. For each class, the class name and its closely related categories (e.g., cabinet, cupboard, bureau, china cabinet) are individually inserted into these templates, tokenized [97], and processed through a CLIP text encoder. Importantly, each synonym generates its own unique embedding, allowing the model to capture subtle semantic differences. However, during the classification process, these separate embeddings are all associated with a single, general class name (e.g., ‘cabinet’) for output, which aids in LLM room classification and scene graph generation. This approach allows for on-the-fly computation of embeddings for new objects without additional training and enhances the model’s robustness to different object naming conventions.
The classification process involves both out-of-vocabulary and in-vocabulary classifiers. The out-of-vocabulary classifier, which applies mask pooling to the frozen CLIP backbone’s feature maps, is designed to maintain CLIP’s open-vocabulary recognition ability for unseen classes. The in-vocabulary classifier, which pools the pixel decoder’s output feature maps, is optimized for better performance on classes seen during training. Both compute classification scores using cosine similarity between the pooled features and each of the synonym embeddings, followed by softmax with a temperature of 0.07. The highest score among synonyms determines the class prediction. These scores from both classifiers are then combined using geometric ensembling, leveraging the strengths of both classifiers to improve overall performance on seen and unseen classes. Importantly, experiments have shown that using both classifiers also improves performance in out-of-vocabulary classes, demonstrating the synergistic effect of this approach. Candidate masks are filtered out based on classification scores or if they belong to the background class, with further filtering out of ambiguous or inconsistent masks where the percentage of pixels belonging to majority classes falls below a specified threshold. Remaining masks $\mathcal{M}_{o}$ are associated with their corresponding object class $o$, using the general class name. Objects like ceilings, floors, and walls are segmented but excluded from object mapping to improve robustness against outlier masks. The classes for the classifiers consist of classes from COCO Panoptic dataset [98, 99, 100] (used to train FC-CLIP) and additional class labels relevant to indoor environments, along with their synonyms. All class labels are mapped to appropriate general class names, with provisions for adding new classes as needed. This process enables FC-CLIP to perform effective zero-shot segmentation across a wide range of object classes, including unseen ones while adapting to various naming conventions and ensuring consistent class names for semantic object mapping.
@@ IV-A2 Semantic Object Mapping
The 2D segmentation mask, $\mathcal{M}_{o}$ (for object $o$), obtained from the model described in Section IV-A1 are used to estimate the object’s center $(u_{o},v_{o})$ in the camera plane. These center points are then projected from the camera plane to 3D coordinates $P_{o}^{c}=(X_{o}^{c},Y_{o}^{c},Z_{o}^{c})$ in the camera frame, using depth frame from RGB-D camera, and camera intrinsics. This approach was selected over alternatives such as using 3D bounding boxes or computing the mean of the point cloud, considering simplicity and computational complexity. These object observations are subsequently converted from the camera frame to the world frame using the pose from the robot’s odometry. We employ an off-the-shelf SLAM toolbox [101] that utilizes the robot’s onboard LiDAR and legged-inertial odometry to build a 2D map of the environment as well as estimate the pose of the robot.
@@
The detected objects ($P_{o}^{c}$ in the camera frame) are transformed to the world frame $P_{o}^{w}$ using the estimated pose and associated with previous map objects. The map objects are observed in prior frames and frame objects are objects observed in the current frame. All frame objects observed in the first frame are considered as map objects and if there are multiple map objects belonging to the same class, they are given a different instance id. The instance id is the instance count of a particular class in the scene and is thus initialized at 1 and incremented with new observations. For the subsequent frames, the frame object and map object are associated using density-based clustering (DB-SCAN) [102] based assignment as it offers a good balance between accuracy and computational speed compared to more complex methods like Hungarian matching [103]. Frame objects are projected into the world coordinate frame and DB-SCAN is applied to group frame and map objects of the same class, using a local radius threshold of $\epsilon$ meters. For each resulting cluster, only one map object is retained, while the others are discarded. DBSCAN is effectively able to handle grouping without using heuristics, such as when a large object is detected as several smaller bounding boxes across consecutive frames.
@@
Three cases for grouping objects are considered: for larger objects like tables, doors, and noticeboards, we set $\epsilon=2$ meters to account for their size; for smaller objects, such as books, a reduced threshold of $\epsilon=0.5$ meters is used; and for all other objects, a threshold of $\epsilon=1$ meter is applied. All the objects without any associations are assigned as new map objects with unique instance id, thus forming a semantic object map $\mathcal{S}$. The system handles updates in object locations following loop closure and local bundle adjustment using a pose graph-based approach. Each object in the semantic object map $\mathcal{S}$ is associated with the keyframe in which it was first observed. A relative transform $T_{o}^{k}$ between the object and its associated keyframe in the pose graph is maintained. When loop closure or local bundle adjustment occurs and keyframe poses are updated, the location of each object is adjusted by applying the updated keyframe pose to the stored relative transform: $T_{o}^{w}=T_{k}^{w}\cdot T_{o}^{k}$, where $T_{o}^{w}$ is the updated world pose of the object, $T_{k}^{w}$ is the updated pose of the keyframe in the world coordinates, and $T_{o}^{k}$ is the relative transform between the object and its keyframe. An alternative approach involving bundle adjustment using all keyframes that observe an object was considered but not chosen due to its higher complexity and computational cost, favoring simplicity and efficiency in our current system. Our method ensures that object locations remain consistent with the refined map and robot trajectory, maintaining the integrity of the semantic object map even as the underlying SLAM system refines its estimates.
@@ IV-A3 Hierarchical Scene Graph
Indoor environments are inherently structured into distinct semantic components, primarily floors and rooms. These spaces are typically categorized using familiar natural language terms such as ‘office,’ ‘kitchen,’ and ‘bedroom,’ based on their intended function and the objects they contain. The consistency of these labels across various indoor settings makes them highly informative semantic descriptors. For instance, a room labeled ‘kitchen’ has a significantly higher likelihood of containing a microwave compared to one designated as an ‘office’. This semantic structuring of indoor spaces lends itself to efficient representation through 3D scene graphs with room hierarchies, as proposed in prior works [46, 47, 15, 16]. Such representations not only capture the physical layout of indoor environments but also encode the semantic relationships between spaces and objects, facilitating more intuitive and context-aware spatial understanding and navigation. The indoor spaces consist of several rooms connected by walkways called corridors. Hence, it is natural to find multiple clusters of objects while exploring such environments. Making use of this observed property, we propose classifying the scene into multiple areas by applying density-based clustering (DBSCAN [102]) on the coordinates of the map objects. For each identified cluster, we leverage LLMs to determine the most appropriate room label. Our approach involves inputting the labels of all objects present in each area cluster to the LLM, along with a curated list of candidate room labels and general guidelines to ensure consistency across the environment. The generation of these labels occurs in real-time and on-device, utilizing a memory-efficient 4-bit quantized version of the LLaMA 3 [104] model. While prior works rely on language-aligned visual embedding models to find the best match [16, 15, 20], our experiments indicate that modeling the influence of certain combinations of objects toward accurate room classification is fairly difficult. Hence, we leverage an LLM to identify labels that are consistent with the general notion of such semantic concepts.
The efficient storage of the scene graph is achieved through a tree structure. Each node in the tree contains properties such as its hierarchy level, text label, class-specific unique ID relative to the parent node, location in the map, and parent and child relationships as shown in Fig. 2. This structure is easily converted to a Python dictionary format, allowing direct input as a prompt to the LLM-based planner, as detailed in Section IV-B. Our approach combines density-based clustering for spatial organization, LLM-powered semantic labeling for accurate room classification, and an efficient tree-based representation for the scene graph. This integrated method offers a robust and flexible approach to understanding and representing complex indoor environments, overcoming the challenges of traditional visual embedding models in capturing the nuanced relationships between objects and room types. By leveraging the semantic understanding capabilities of LLMs, we achieve a more intuitive and context-aware representation of indoor spaces, which can significantly enhances the agent’s ability to navigate, localize objects, and comprehend the overall layout of the indoor scene for autonomous task execution.
@@ IV-B LLM Planner
The LLM planner leverages GPT-4-Turbo [41] to navigate and interact with the environment through three core action primitives: goto(<room name> or <object name>), search_room(<room name>), and explore_globally(). The goto command directs the robot to the specified room or object and queries a visual agent to verify task completion. The search_room command moves the robot to the specified room and then rotates to detect nearby objects. The explore_globally command initiates frontier-based exploration of the environment. The documentation of these commands with detailed explanations and use-case scenarios, is provided to the LLM (see Fig. 4 for prompting details). The LLM’s grounding in the environment is achieved through a hierarchical scene graph formatted as JSON, which stores information about rooms and the objects within them (as described in Section IV-A3). Leveraging this scene graph, the LLM selects an appropriate action and fills in the necessary arguments to accomplish the user-defined task, while also providing reasoning for its decision. The resulting output is directly executed as a function call to the robot API, implemented using the ‘Safe Robot Controller’ described in Section IV-C.
@@
12:           task_accomplished $\leftarrow$ ask_agent(‘Is task done?’)
@@
The plan generation framework initiates with a 360 environmental scan if the scene graph is empty. The LLM then generates an action based on the given task and hierarchical scene graph. Following action execution, if the task remains incomplete, the LLM receives an updated scene graph and generates a subsequent action. Error scenarios—stemming from incorrect syntax, low-level controller issues, or environmental constraints like inaccessible target points—trigger feedback messages explaining the error, prompting the LLM to produce alternative actions. After the initial action, a command history is maintained to prevent action repetition. This iterative action generation process is formalized in Algorithm 1. The action primitives used by the LLM are elaborated as follows:
@@
The goto primitive enables the robot to navigate to a specified room or approach a particular object within the environment. When tasks involve navigating to an object, the LLM planner is instructed via system prompt to provide both the object’s name and the room it resides in as arguments. For example, to locate a ‘printer in the office,’ the command would be goto(office-2, printer-1). Conversely, for tasks requiring navigation to a specific room, only the room name is specified, such as goto(office-1) for ‘an office with three computers.’ In each command, the name consists of a label and a unique instance ID, where object labels are assigned during semantic object mapping, and instance IDs and room labels are generated during hierarchical scene graph creation. This primitive also supports task completion verification by interacting with a visual agent upon arrival at the destination.
@@
This primitive commands the robot to navigate to the specified room and perform a 360 rotation within it. This action is designed to explore a partially observed room, enabling the robot to detect and recognize objects in the scene and update its scene graph. Even if the robot has previously visited the room, limitations in perception or environmental changes may have caused it to miss or misidentify objects. By re-observing the room, the robot gathers more complete and up-to-date information, which is critical for accurate planning and task execution. While this command might be redundant in simulated environments with ground truth perception and segmentation, it is essential in real-world settings where the processed data from the sensors can be imperfect or incomplete. The ability to update the scene graph with newly observed objects as well as updated locations ensures that the LLM has access to the most current environmental information for decision-making.
@@
When the scene graph is empty or lacks sufficient information for the LLM to utilize the other commands effectively, the explore_globally() primitive is employed. This action initiates an autonomous exploration phase, typically lasting one minute, using a frontier-based exploration strategy as detailed in Section IV-C3 to enable identification of unexplored areas. The robot continuously updates its scene graph using new observations from onboard sensors during the exploration phase. Previously unobserved aspects of its surroundings are updated by rotating the quadruped around its vertical axis. The enhanced scene representation enables the LLM planner to generate more specific and task-relevant commands. The duration of the exploration is selected to balance thoroughness with operational efficiency, ensuring that the robot gathers enough information to proceed with task-specific actions without unnecessary delay. This primitive is vital when the robot starts in an unknown environment or when significant portions of the environment remain unexplored, providing the foundational knowledge required for effective planning and execution of subsequent tasks.
IV-C Safe Robot Controller